OpenStudy (anonymous):

Derivative. variable in base and exponent. f(x)=x^{2x+3} I haven't been able to find the steps to solve a derivative with x in the base and the exponent. How do I solve this type of derivative?

OpenStudy (anonymous):

\[f(x)=x ^{2x+3}\] (formatted)

OpenStudy (dumbcow):

best way is to use logs \[y = x^{2x+3}\] \[\ln y = (2x+3) \ln x\] now differentiate \[\frac{1}{y} \frac{dy}{dx} = \frac{2x+3}{x} +2\ln x\] \[\frac{dy}{dx} = x^{2x+3}(\frac{2x+3}{x} +2\ln x)\]

OpenStudy (anonymous):

so rewrite it as "y=", take ln, then use implicit differentiation?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

I see. dy/dx trips me up sometimes, so I stick to y' and looks like I get the same answer. \[y=x^{2x+3}\] \[\ln (y) = (2x+3) \ln(x)\] then take derivative of both sides. \[\frac{ y' }{ y }=2\ln(x)+(2x+3)\frac{ 1 }{ x }\] \[y'=y(2\ln(x)+\frac{ 2x+3 }{ x })\] \[y'=x^{2x+3}(2\ln(x)+\frac{ 2x+3 }{ x })\] Thanks!

OpenStudy (dumbcow):

yw